1. Factor completely x2 ax +bx + ab 4x2-49y2 2. Factor
Answer #1

1. x-tax +bx+aь . (x2+ax) + (Lx+42) (х+а) +ь(х+4) = (а +а) (+ь) 2. 4- 499 = 2*x*- 7** = (21) -- 1 a-b= (a-)(a+) = (2x-132 (24